# A Comprehensive Evaluation of Parameter-Efficient Fine-Tuning on Method-Level Code Smell Detection

### v1

@@ 3.3.1. Selected Code Smells
In this study, we intend to use LLMs to detect two types of method-level code smells, i.e., Complex Conditional and Complex Method. The definitions of the two types of code smells are provided below:
  • Complex Conditional (CC): This code smell occurs when a conditional statement is excessively intricate (sharma2017house, ).
  • Complex Method (CM): This code smell occurs when a method exhibits high cyclomatic complexity (sharma2017house, ).
We chose CC and CM because they have been widely studied in research on DL-based code smell detection (e.g., (sharma2021code, ; ho2023fusion, ; zhang2024code, ; alazba2023deep, )). Another reason for selecting CC and CM is the availability of detection tools, such as DesigniteJava (designitejava, ), which can assist us in collecting candidate code samples that may exhibit these two types of code smells. We excluded very simple smells (e.g., Long Method), which can be easily identified by basic techniques. In contrast, detecting the two selected method-level code smells requires LLMs to spot specific features, pay attention to cohesion and coupling aspects, and analyze the entire method for its structural properties (sharma2021code, ), thereby expanding the ambition and enhance the applicability of this study.
@@ 3.3.2. Java Repository Selection
We used GitHub REST API [^1] https://docs.github.com/en/rest?apiVersion=2022-11-28 to search Java repositories from GitHub. From the extensive array of available Java projects hosted on GitHub, we set the following criteria:
  • The repositories should be created between January 1, 2020 and April 30, 2024.
@@
These criteria ensured that the selected Java repositories were recent, popular, active, and non-trivial Java repositories. By employing the stringent filtering criteria, we obtained a list of 294 GitHub repositories coded in Java.
@@ 3.3.3. Potential Code Smell Detection
DesigniteJava is an open-source tool that supports detection of 18 design and 10 implementation smells, which have been extensively used in many studies as a mature code smell detector (sharma2021qscored, ; shah2023mining, ; barbosa2020revealing, ). In this paper, we had developed DesigniteJava to detect potential CC and CM in the Java repositories we collected in Section 3.3.2. After this step, we got a list of code smells, consisting of the project name, package name, class name, and method name where each detected code smell occurs. It is important to note that the code smell detection results from DesigniteJava may not be entirely accurate, as the detection results can contain both false positives and false negatives. False positives occur when the tool incorrectly flags code as exhibiting CC or CM, while false negatives happen when actual code smells are missed. These inaccuracies stem from the rule-based and heuristic nature of DesigniteJava, which relies on predefined criteria that may not capture all nuances of the code.
@@ 3.3.4. Data Preparation
CodeSplitJava (codesplitjava, ) is a utility program designed to partition methods from Java source code into distinct files. We used CodeSplitJava to parse the 294 Java repositories collected in Section 3.3.2 into code files. Each of these code files corresponds to a method in the 294 repositories, and its path reflects the hierarchical structure of the method (i.e., namespaces and packages are converted into folders). By correlating the list of detected code smells collected in Section 3.3.3, with the paths of the code files generated by CodeSplitJava, we identified the method-level code snippets with code smells and those without (i.e., Clean Code).
We removed any duplicate samples within each type of code smell to ensure uniqueness. Additionally, we limited the maximum token length of the samples to 1,024, filtering out any that exceeded this limit. Table 1 provides the number of potential positive samples after each step for the two types of code smells. In total, we collected 4,266 potential positive code samples containing CC and 6,479 containing CM. For both CC and CM, we introduced an equal number of potential negative samples (i.e., code samples without code smells as identified by DesigniteJava) for the subsequent steps. We also verified that all negative samples adhered to the maximum token limit and were free of duplicates. However, the quality of these datasets may be compromised as they could contain false positives and false negatives, requiring further refinement.
@@
Complex Conditional 19,754 4,589 4,266
Complex Method 14,979 9,781 6,479
3.3.5. Manual Validation
After the above steps, we obtained an equal number of potential positive and negative samples for both CC ($4,266\times 2=8,532$) and CM ($6,479\times 2=12,958$). To ensure the quality of the datasets for these two types of code smells and to exclude false positives and false negatives, a manual validation was conducted by the first and fourth authors. Initially, a pilot validation was performed at the 95% confidence level, with a margin of error of less than 5%, resulting in the random selection of 368 samples from CC and 374 from CM. The first and the fourth author annotated the selected samples independently, and the pilot annotations were subsequently compared. Cohen’s Kappa coefficient (jacob1960coefficient, ) was used to measure the inter-rater agreement for both CC and CM. The resulting Kappa coefficient for CC was 0.85, while for CM it was 0.81, both exceeding 0.7, indicating a strong level of agreement between the two annotators. Following the pilot validation, the two authors manually annotated the remaining code samples. Throughout the entire validation process, including the pilot phase, any disagreements between the first and fourth authors were resolved through discussions involving the second author, ensuring consistency in the annotations. Overall, we identified 2,708 false positives and 29 false negatives for CC, and 2,376 false positives and 20 false negatives for CM. The numbers of both positive and negative samples for CC and CM after manual validation are presented in Table 2.
Table 2. Number of positive and negative samples after data annotation
@@ 3.3.5. Manual Validation
Complex Conditional 1,587 6,945 8,532
Complex Method 4,123 8,835 12,958
@@ 3.3.6. Data Split
After manually reviewing the entire dataset, we split the code samples for both CC and CM into training, validation, and test sets following an 8:1:1 ratio, as employed in previous studies (yang2024federated, ; liu2023benchmarking, ). Specifically, the positive and negative samples for each method-level code smell were distributed according to the same 8:1:1 proportion. This approach ensures a consistent distribution of positive and negative cases, providing a solid foundation for model training, validation, and evaluation. The number of code samples in the datasets of CC and CM after this step is presented in Table 3. We used these clean and high-quality datasets as our final datasets for a series of experiments.
@@
Set Complex Conditional Complex Method
#Positive #Negative #Total #Positive #Negative #Total
Train 1,269 5,557 6,826 3,299 7,067 10,366
Valid 159 694 853 412 884 1,296
Test 159 694 853 412 884 1,296
Total 1,587 6,945 8,532 4,123 8,835 12,958
@@ 3.4. Base Model Selection
In this work, we aim to explore the effectiveness of PEFT methods on method-level code smell detection task. To ensure a thorough analysis, we have chosen to use both small LMs and LLMs as our base models. Note that we deem models with 1B or more parameters as LLMs, while referring to those with fewer parameters as small language models.
First, the selected LMs should be open-source as PEFT methods require access to their parameters. Second, the models should support classification tasks as we frame code smell detection as a binary classification problem. Third, the base models used in this study must be able to be fine-tuned and tested on a single 48GB GPU (our maximum computational resource) without encountering memory overflow. Fourth, the selected models should either be popular small LMs or recent advanced LLMs to ensure they represent state-of-the-art capabilities.
Based on the above criteria, we employed 4 small LMs (i.e., CodeBERT (feng2020codebert, ), GraphCodeBERT (guo2021graphcodebert, ), UnixCoder (guo2022unixcoder, ), and CodeT5 (wang2021codet5, )) and 6 LLMs (DeepSeek-Coder-1.3B (guo2024deepseekcoder, ), DeepSeek-Coder-6.7B (guo2024deepseekcoder, ), StarCoderBase-1B (li2023starcoder, ), StarCoderBase-3B (li2023starcoder, ), StarCoderBase-7B (li2023starcoder, ), and CodeLlama-7B (roziere2024codellama, )) for our experiments. These models feature a range of transformer architectures, including encoder-only, encoder-decoder, and decoder-only modes. All the models we selected have demonstrated encouraging results across various code understanding tasks like code clone detection, defect detection, etc. (liu2024vul, ; karmakar2021what, ; lu2021codexglue, ; li2022automating, ; tang2023csgvd, ; zeng2022an, ; chen2023diversevul, ; wang2023one, ; fu2022vulrepair, ; nguyen2024empirical, ; su2024distilled, ). Table 4 presents the details of the 10 base models, including their parameter sizes and architecture.
@@
Small LMs CodeBERT 125M Encoder-only
@@
LLMs DeepSeek-Coder-1.3B 1.3B Decoder-only
@@ 3.5. Metrics
We use metrics that align with previous research (sharma2021code, ; ho2023fusion, ; pecorelli2019comparing, ) to maintain consistency in evaluating the effectiveness of PEFT methods for code smell detection on LLMs. Specifically, we utilize precision, recall, F1-score, and MCC (Matthews Correlation Coefficient) for assessing the performance of detecting CC and CM. We include MCC because, unlike other metrics like F1-score, accounts for true negative instances (yao2020assessing, ). These metrics are calculated using the following formulas:
@@ 3.6. Implementation Details
The code for our experiments is implemented by the PEFT library [^2] https://huggingface.co/docs/peft/index from Hugging Face. After reviewing prior studies (liu2023empirical, ; liu2024delving, ; li2024comprehensive, ; wang2023one, ; weyssow2023exploring, ) and observing the results of pilot experiments, we set the learning rate to $3e^{-4}$ for all PEFT methods and $1e^{-5}$ for full fine-tuning the small LMs. Throughout all the fine-tuning experiments, we utilized the AdamW optimizer (loshchilov2017decoupled, ) and fixed the Transformer input window size at 1,024. All the models were trained for 10 epochs, and we selected the epochs at which models exhibit minimum loss on the validation sets for inference. We executed all models on a single Linux GPU server equipped with a 48G NVIDIA RTX 6000 Ada Generation GPU. Additionally, we have released our datasets and source code online for replication purpose (replicationPackage, ) where more details, including hyper-parameter values, can be found.

### v2

@@ 3.3.1. Selected Code Smells
In this study, we intend to use LMs with PEFT methods to detect four types of code smells, comprising three at the method level (i.e., Complex Conditional, Complex Method, and Feature Envy) and one at the class level (i.e., Data Class). The definitions of these four code smell types are provided below:
We selected CC, CM, FE, and DC because they are among the most extensively studied code smells in prior research on ML- and DL-based code smell detection (e.g., (sharma2021code, ; ho2023fusion, ; zhang2024code, ; alazba2023deep, ; fontana2016comparing, ; barbez2020machine, ; kim2017finding, ; liu2023deep, ; liu2018deep, ; skipina2024automatic, ; yang2024feature, ; fontana2013code, ; madeyski2020mlcq, )). Moreover, these code smells are frequently observed in real-world software projects and are known to significantly impact code quality, maintainability, and the likelihood of introducing defects (fowler1999refactoring, ). For example, CC and CM can lead to convoluted control flows that hinder code readability and increase cognitive load on developers, while FE and DC often indicate bad modularization and poor class design, which complicate maintenance and evolution (fowler1999refactoring, ; sjoberg2012questioning, ; khomh2009exploratory, ).
Another reason for choosing these specific smell types is the availability of detection tools, such as DesigniteJava (designitejava, ) and PMD (pmd, ), which facilitates the collection of candidate code samples potentially exhibiting these smells. We excluded overly simple smells (e.g., Long Method) that can be easily identified using basic heuristics. In contrast, detecting the four selected code smells requires LLMs to capture subtle features, consider aspects such as cohesion and coupling, and analyze the overall structure of a code sample (sharma2021code, ), which expands the ambition and enhances the applicability of this study.
At the same time, identifying these selected code smells allows us to assess the effectiveness of LMs fine-tuned with PEFT methods in both method-level and class-level code understanding. CC, CM, and FE are generally considered method-level code smells, while DC represents a class-level code smell. However, detecting FE requires comprehension of not only the focal method but also its surrounding class context and inter-class dependencies (fowler1999refactoring, ; tsantalis2009indentification, ; liu2023deep, ; alkharabsheh2021analysing, ). Therefore, it also demands class-level code understanding. The selection of these four code smells allows our study to address a key gap in existing research related to PEFT methods (li2024comprehensive, ; liu2023empirical, ; wang2023one, ; liu2024delving, ; weyssow2023exploring, ), which has primarily concentrated on the effectiveness of PEFT techniques on method-level code-related tasks.
@@ 3.3.2. Java Repository Selection
To construct a code smell dataset that accurately represents modern coding practices, we selected Java repositories created in the past six years from GitHub. We used the GitHub REST API [^3] https://docs.github.com/en/rest?apiVersion=2022-11-28 to gather available Java projects that meet the following criteria:
  • The repositories should be created between January 1, 2019 and April 30, 2024.
@@
These criteria ensured that the selected Java repositories were recent, popular, active, and non-trivial. By employing these stringent filtering criteria, we obtained a list of 401 GitHub repositories coded in Java.
@@ 3.3.3. Potential Code Smell Detection
DesigniteJava is an open-source tool capable of detecting 18 design smells and 10 implementation smells, whereas PMD is a rule-based static analysis tool that identifies common programming issues, including code smells. Both of them have been extensively used in prior studies (sharma2021qscored, ; shah2023mining, ; barbosa2020revealing, ; fernandes2016review, ; aniche2018code, ; wu2024ismell, ). In this paper, we leveraged DesigniteJava to detect potential instances of CC, CM, and FE, and used PMD to detect potential instances of DC. The output of this step is a list of detected code smells, each annotated with the corresponding project, package, class, and, where applicable, method name. It is worth noting that the code smell detection results from DesigniteJava and PMD may not be entirely accurate, as the detection results can contain both false positives and false negatives. False positives occur when the tool incorrectly flags code as exhibiting certain type of code smell, while false negatives happen when actual code smells are missed. These inaccuracies stem from the rule-based and heuristic nature of DesigniteJava and PMD, which relies on predefined criteria that may not capture all nuances of the code.
@@ 3.3.4. Data Preparation
CodeSplitJava (codesplitjava, ) is a utility program designed to extract methods or classes from Java source code and organize them into separate files. We employed CodeSplitJava to parse the 401 Java repositories collected in Section 3.3.2, generating individual code files for each method or class. The file paths reflect the hierarchical structure of the original code (i.e., namespaces and packages are converted into nested folders). By aligning the list of detected code smells obtained in Section 3.3.3 with the paths of the code files generated by CodeSplitJava, we were able to collect code snippets that were labeled by heuristic-based detectors as either containing code smells or being free of them.
@@
Complex Conditional 23,984 7,768 6,901
Complex Method 20,171 14,687 9,819
Feature Envy 2,152 1,113 337
Data Class 9,672 8,501 7,121
We removed any duplicate samples within each type of code smell to ensure uniqueness. Additionally, we limited the maximum token length of the samples to 1,024, filtering out any that exceeded this limit. This decision was made to accommodate the input constraints of the models used in our experiments (see Section 3.6). Specifically, the input window size for the LMs was set to 1,024 tokens to align with GPU memory limitations. Samples exceeding 1,024 tokens undergo truncation, which may result in the loss of critical contextual information. For instance, key sections of code relevant to the detection of code smells could be omitted, potentially impacting the model’s performance. Therefore, we set a hard limit of 1,024 tokens to avoid such issues, ensuring that all samples in our dataset can be fully utilized by the models without sacrificing the integrity of their content. Table 1 provides the number of potential positive samples after each step for the four types of code smells. Note that although FE is a method-level code smell, we retained the entire class that contains the method as the sample. This decision aligns with the definition of FE (see Section 3.3.1), which emphasizes the method’s interaction with its enclosing class. Providing the full class context ensures that the code sample contains sufficient structural and semantic information for FE detection.
3.3.5. Manual Review
After the above steps, we obtained two sets of samples for each code smell: a potential positive set and a potential negative set, with the latter being significantly larger in size. To construct a high-quality and balanced dataset, the first and the fourth authors carried out a two-stage manual review process, as illustrated in Figure 4.
In the first stage, we reviewed all potential positive samples manually. Each sample was carefully examined to identify and eliminate false positives. If a sample in the potential positive set had been incorrectly labeled as positive by DesigniteJava or PMD, it was reclassified as negative and moved to the negative set; otherwise, it was transferred into the positive set. This stage yielded a verified positive set and a smaller verified negative set. In the second stage, in order to enlarge the negative set to achieve dataset balance, we randomly selected samples from the potential negative set for manual validation. Similar to the first stage, samples confirmed as true negatives were added to the negative set, while those identified as false negatives were reclassified into the positive set. This process was repeated iteratively until an equal number of manually reviewed positive and negative samples was achieved. During this stage, all randomly selected potential negative samples were required to be unique, adhere to the predefined maximum token limit, and be representative. To ensure the representativeness of negative samples across the four types of code smells, we established the following criteria: for CC, negative samples must contain at least one conditional statement; for CM, negative samples must contain at least 15 lines of code; and for FE and DC, negative samples must contain at least 30 lines of code, as we extract the corresponding class for these two smells. These constraints help filter out trivial or uninformative code snippets and ensure that the negative samples are sufficiently complex and realistic, thereby contributing to a more robust and reliable dataset. Through this two-stage review, we ensured that the final dataset for each code smell type was both balanced and of high quality.
Figure 4. The process of two-stage manual review in dataset construction
Table 2. Number of positive and negative samples after manual review
@@ 3.3.5. Manual Review
Complex Conditional 4,604 4,604 9,208
Complex Method 5,748 5,748 11,496
Feature Envy 307 307 614
Data Class 6,922 6,922 13,844
In order to ensure the consistency of data annotation, a pilot review was performed by the first and the fourth authors before conducting the formal manual review process. For each code smell type (i.e., CC, CM, FE, and DC), 30 potential positive samples and 30 potential negative samples labeled by heuristic-based tools were randomly selected. The two authors independently examined these samples and then compared their labeling results. Cohen’s Kappa coefficient (jacob1960coefficient, ) was used to assess inter-rater agreement between the two authors. The Kappa coefficient for CC was 0.83, for CM was 0.80, for FE was 0.77, and for DC was 0.87. All values exceeded the threshold of 0.7, indicating a strong level of agreement between the annotators. Following the pilot review, the two authors held a discussion to reconcile discrepancies and formalize the labeling criteria below:
  • CC: Labeled as positive if the method contains at least one conditional statement that either (i) has a nesting depth of three or more, making the method hard to read, or (ii) includes compound boolean expressions involving three or more operands, exhibiting complex logic that makes the condition challenging to evaluate and understand.
  • CM: Labeled as positive if the method simultaneously satisfies both (i) three or more execution paths with complex control flow, and (ii) an implementation that is difficult to comprehend or follow.
  • FE: Labeled as positive if the method inside the input class satisfies both (i) over 50% of its data access operations, such as method calls and field accesses, target members of external classes rather than its own class, and (ii) relocating the method to one of these external classes would improve encapsulation and cohesion.
  • DC: Labeled as positive if the class contains (i) primarily fields accompanied by trivial getters/setters (at least 80% of its fields), and (ii) lacks methods that implement significant logic or domain behavior, indicating its role as a passive data holder.
These criteria were grounded in established definitions from prior studies (sharma2017house, ; fowler1999refactoring, ; fontana2016comparing, ; pietrzak2006leveraging, ) and refined based on insights from our pilot study. During the formal manual review, the two authors collaboratively labeled all samples (i.e., each sample was assigned to one of the two authors), which took approximately 300 person-hours in total. Whenever questions or ambiguities arose, the second author, who is an expert in code smell detection, was consulted. In such cases, all three authors discussed the sample until consensus was reached, ensuring annotation reliability and consistency. The final count of positive and negative samples, following the manual review, is presented in Table 2.
@@ 3.3.6. Data Split
In this step, we split the code samples of CC, CM, FE, and DC into training, validation, and test sets following an 8:1:1 ratio, as employed in previous studies (yang2024federated, ; liu2023benchmarking, ). For each type of code smell, we ensured that the number of positive and negative samples remains equal across the training, validation, and test sets, thereby maintaining class balance throughout the entire dataset. Table 3 outlines the number of code samples for each code smell type after the data split. We then used these clean and high-quality datasets as our final datasets for a series of experiments.
@@
Code Smell Train Valid Test Total
Complex Conditional #Positive 3,684 460 460 4,604
#Negative 3,684 460 460 4,604
Complex Method #Positive 4,600 574 574 5,748
#Negative 4,600 574 574 5,748
Feature Envy #Positive 245 31 31 307
#Negative 245 31 31 307
Data Class #Positive 5,538 692 692 6,922
#Negative 5,538 692 692 6,922
@@ 3.4. Base Model Selection
In this work, we aim to explore the effectiveness of PEFT methods for the task of code smell detection. To ensure a thorough analysis, we included both small and large LMs as our base models. Consistent with the definition adopted by Weyssow et al. (weyssow2023exploring, ), we classified models with 1B or more parameters as LLMs, and those with fewer parameters as SLMs.
@@
SLMs CodeBERT 125M Encoder-only
@@
LLMs StarCoderBase-1B 1.1B Decoder-only
StarCoderBase-3B 3B
DeepSeek-Coder-1.3B 1.3B
@@
First, the selected LMs should be open-source as PEFT methods require access to their parameters. Second, the models should support classification tasks as we frame code smell detection as a binary classification problem. Third, the base models used in this study must be able to be fine-tuned and tested on a single 40GB GPU (our maximum computational resource) without encountering memory overflow. Fourth, the selected models should either be popular SLMs or recent advanced LLMs to ensure they represent state-of-the-art capabilities.
Based on the above criteria, we employed 4 SLMs (i.e., CodeBERT (feng2020codebert, ), GraphCodeBERT (guo2021graphcodebert, ), UnixCoder (guo2022unixcoder, ), and CodeT5 (wang2021codet5, )) and 5 LLMs (StarCoderBase-1B (li2023starcoder, ), StarCoderBase-3B (li2023starcoder, ), DeepSeek-Coder-1.3B (guo2024deepseekcoder, ), DeepSeek-Coder-6.7B (guo2024deepseekcoder, ), and CodeLlama-7B (roziere2024codellama, )) for our experiments. To ensure a comprehensive analysis, we deliberately included models representing all three major Transformer architectures: encoder-only, encoder-decoder, and decoder-only. Although decoder-only models were originally designed for generative tasks (vaswani2017attention, ; radford2019language, ), prior studies have shown their competitive performance in sequence classification tasks as well (wolf2020transformers, ; kariyappa2024progressive, ). Therefore, we also included decoder-only models as base models in our study. Additionally, all the models we selected have demonstrated encouraging results across various code understanding tasks like code clone detection, defect detection, etc. (liu2024vul, ; karmakar2021what, ; lu2021codexglue, ; li2022automating, ; tang2023csgvd, ; zeng2022an, ; chen2023diversevul, ; wang2023one, ; fu2022vulrepair, ; nguyen2024empirical, ; su2024distilled, ). Table 4 presents the details of the 9 base models, including their parameter sizes and architectural types.
@@ 3.5. Metrics
We use metrics that align with previous research (sharma2021code, ; ho2023fusion, ; pecorelli2019comparing, ) to maintain consistency in evaluating the effectiveness of PEFT methods for code smell detection on LLMs. Specifically, we utilize precision, recall, F1-score, and MCC (Matthews Correlation Coefficient) for assessing the performance of detecting CC and CM. We include MCC because, unlike other metrics like F1-score, MCC accounts for true negative instances (yao2020assessing, ). These metrics are calculated using the following formulas:
@@ 3.6. Implementation Details
The code for our experiments is implemented by the PEFT library [^4] https://huggingface.co/docs/peft/index from Hugging Face. After reviewing prior studies (liu2023empirical, ; liu2024delving, ; li2024comprehensive, ; wang2023one, ; weyssow2023exploring, ) and observing the results of pilot experiments, we set the learning rate to $3e^{-4}$ for all PEFT methods and $1e^{-5}$ for full fine-tuning the SLMs. Throughout all the fine-tuning experiments, we utilized the AdamW optimizer (loshchilov2017decoupled, ) and fixed the Transformer input window size at 1,024. The choice to limit the input length to 1,024 tokens was driven by hardware constraints. As the GPU memory consumption of transformer-based models grows roughly quadratically with input length (tay2022efficient, ), this setting enables all selected base models to train within available GPU memory capacity. All the models were trained for 10 epochs, and we selected the epochs at which models exhibit minimum loss on the validation sets for inference.
Table 5 shows the hyperparameter settings of different PEFT methods. We adopted a batch size of 1 for all experiments to avoid GPU out-of-memory issues. We used a fixed random seed (42) across all runs to ensure experimental reproducibility. For prompt tuning and prefix tuning, we used 20 virtual tokens, which are learnable vectors prepended to the input and optimized during training. This setting is widely used in the literature and offers a balance between representational capacity and computational cost (li2021prefix, ; weyssow2023exploring, ). As for LoRA, we set the rank $r$ to 8, the scaling factor $\alpha$ to 16, and the dropout rate to 0.1 to prevent overfitting, following configurations reported in (liu2024delving, ). For both LoRA and ${IA}^{3}$, we applied parameter-efficient adaptation to the core components of the Transformer architecture, including the query ($Q$), key ($K$), and value ($V$) projections in the attention layers. In the case of $(\text{IA})^{3}$, we also included the Feed-Forward Network (FFN) for adaptation (see Section 2.3). We executed all models on a single Linux GPU server equipped with a 40GB NVIDIA A100 GPU. Additionally, we have released our datasets and source code online for replication purpose (replicationPackage, ) where more details can be found.
Table 5. Hyperparameter settings of different PEFT methods and full fine-tuning
PEFT Method Hyperparameter Setting Batch Size Seed
Prompt Tuning Number of virtual tokens: 20 1 42
Prefix Tuning
LoRA Rank $r$: 8
$alpha$: 16
$dropout$: 0.1
Target modules: $Q$, $K$, and $V$
$(IA)^{3}$ Target modules: $Q$, $K$, $V$, and FFN
Full Fine-tuing None